OpenStudy (anonymous):

simplify ln(sin pi/4) HOWWWW :(

OpenStudy (irishboy123):

start with \(sin \frac{\pi}{4}\) what's that?

OpenStudy (anonymous):

ln[1/√(2)] = ln[ (√2)⁻¹ ] = ln[ (2)^(-1/2) ] =(-1/2)ln2 √2 means 2's 1/2 th power

OpenStudy (anonymous):

@missylulu